Susan deposited 49228 € for 31.3. – 28.8. and received net accrued interest of 322.6 €. The tax at source rate was 30 %. Calculate the nominal annual interest rate (= interest rate which contains both investor's and taxman's shares).


1
Expert's answer
2022-02-07T13:37:21-0500

Days between 31st march to 28th August in a non-reap year=150

Days in the year"=365"


Accrued interest "=\frac{100}{70}×322.6"

"=460.9"


Let r "=" Annual interest rate

"\frac{r}{100}×\frac{1}{365}×150×49228=460.9"


"r=2.28\%\>p.a"